# Latent Assistance Networks: Rediscovering Hyperbolic Tangents in RL

### v1

@@ Abstract
Activation functions are one of the key components of a neural network. The most commonly used activation functions can be classed into the category of continuously differentiable (e.g. tanh) and linear-unit functions (e.g. ReLU), both having their own strengths and drawbacks with respect to downstream performance and representation capacity through learning (e.g. measured by the number of dead neurons and the effective rank). In reinforcement learning, the performance of continuously differentiable activations often falls short as compared to linear-unit functions. From the perspective of the activations in the last hidden layer, this paper provides insights regarding this sub-optimality and explores how activation functions influence the occurrence of dead neurons and the magnitude of the effective rank. Additionally, a novel neural architecture is proposed that leverages the product of independent activation values. In the Atari domain, we show faster learning, a reduction in dead neurons and increased effective rank.
@@ 1 Introduction
Figure 1: Normalized performance training DQN in the Atari domain where the activation of the representation, defined as the final hidden layer, is either ReLU or tanh.
@@
Out of all activation functions, the Rectified Linear Unit (ReLU) (Nair & Hinton, 2010) and its variants (Xu et al., 2015; Klambauer et al., 2017) have emerged as the most widely used and generally best-performing activation functions up until this day (Jarrett et al., 2009; Goodfellow et al., 2016). The strength of the ReLU activation lies in its ability to naturally avoid vanishing gradients when used in deeper networks, in contrast to the continuously differentiable activation functions, such as the sigmoid and the hyperbolic tangent (Glorot & Bengio, 2010).
@@
A common drawback of using the ReLU activation is its limited expressivity in the context of shallow networks (see Figure 2), as well as the phenomenon known as the dying ReLU problem (He et al., 2015; Lu et al., 2019). As training progresses, the number of dying ReLU’s tend to increase, resulting in a dying network (Dubey et al., 2022).
@@
In reinforcement learning (RL) (Sutton & Barto, 2018), even though training results in a large number of dying ReLU’s (Gulcehre et al., 2022; Sokar et al., 2023), the ReLU function remains the most popular activation for performance reasons (Henderson et al., 2018). As a result, continuously differentiable activation functions like the hyperbolic tangent are rarely used (see Fig. 1), even though their symmetrical, bounded shape and smooth gradient landscape has advantages for optimization. Not much attention has been given to investigate and provide a solution to this mysterious lack of performance, despite evidence showing that a hidden layer activated by a hyperbolic tangent is potentially a strong choice, as it leads to a high effective rank and thus network expressivity (Kumar et al., 2021; Gulcehre et al., 2022).
From the perspective of a compressed representation, defined as the activations of the final hidden layer, we therefore investigate the use and improvement of continuously differentiable activations in the reinforcement learning context. This is done by providing an alternative to the conventional way of parameterizing a representation. Specifically, the following contributions are made to the field:
@@
  • We show that, similar to dying ReLU’s, dying hyperbolic tangents are also abundant in the representation when training RL.
  • A Latent Assistance Network (LaN) is introduced, modifying the representation of a network to be the product of two separate, individual activations.
@@
  • We demonstrate that, without tuning any hyperparameters or using any auxiliary losses, the addition of a LaN yields significant performance gains over using a standard continuously differentiable activation in the Atari domain, and reveal how it decreases dying neurons and increases the internal representations’ effective rank.
@@ 2 Preliminaries
Figure 2: A regression of three shallow neural network architectures on a complex sinusoidal function ($y=10*torch.sin(7*x)+15*torch.sin(10*x)+5*torch.cos(5*x)$). The Tanh (LaN) network emerges as the strongest function approximator, even while having less trainable parameters (501 vs 601 for Tanh & ReLU). To make a fair comparison, the Tanh and ReLU networks have one single hidden layer of size 200, while the Tanh (LaN) network has one single hidden layer of size 100. For the Tanh (LaN) network however, we use two parallel linear layers preceding the hidden layer in order to be able to use the single hidden layer as the product of two activations (see Section 3). For experiments comparing deeper networks, we refer the reader to Appendix C.1.
We consider an agent acting within its environment, where the environment is modeled as a discrete Markov Decision Process (MDP) defined as a tuple $(\mathcal{S},\mathcal{A},T,R,\gamma)$. Here, $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the environment’s transition function, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{R}$ is the environment’s reward mapping and $\gamma$ is the discount factor. A replay buffer $B$ is used to store visited states $s_{t}\in\mathcal{S}$ that were followed by actions $a_{t}\in\mathcal{A}$ and resulted in the rewards $r_{t}\in\mathcal{R}$ and the next states $s_{t+1}$. One entry in $B$ contains a tuple of past experience $(s_{t},a_{t},r_{t},s_{t+1})$. The agent’s goal is to learn a policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ that maximizes the expectation of the discounted return $V^{\pi}(s)=\operatorname{\mathbb{E}}_{\tau}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\mid s_{t}=s]$, where $\tau$ is a trajectory following the policy $\pi$.
@@
3 Method
Our method aims at using the symmetrical, stable gradient landscape and the naturally bounded properties of a continuously differentiable activation, while avoiding the pitfalls such as the well-known vanishing gradient problem. To accomplish this, we complement the original linear layer with an additional, independent linear layer. Subsequently, we define the final representation as a product of the two linear layers’ activations. In this section, the continuously differentiable activations as well as the full architecture used in our research are explained.
3.1 Continuously Differentiable Activations
@@
3.2 Latent assistance Network (LaN)
Figure 3: A visualisation of the Latent assistance Network (LaN) architecture combined with the nature DQN architecture applied on a snapshot of the ’SpaceInvaders’ Atari environment (Mnih et al., 2015). A parallel linear layer providing independent hidden layer activations is integrated, where the element-wise product of the two activations represents the final latent representation $z_{t}$ used for downstream learning.
@@ 3.2 Latent assistance Network (LaN)
To mitigate saturating activations, an augmentation of the conventional representation architecture is proposed. The networks’ original hidden layer is defined as $z^{enc}(x)=f(A_{1}x+B_{1})$, with $A_{1}$ and $B_{1}$ representing the layer parameters, and with the function $f()$ that represents a nonlinear activation function while $x$ is the set of activations from the previous layer. In order to reduce the number of neurons that are fully saturated (same value for all inputs, similar to collapse), the Latent assistance Network augments the original representation by a parallel representation layer $z^{*}$ that can be seen as a single highway layer with a closed carry gate (Srivastava et al., 2015). The final representation is defined as the element-wise product between the aforementioned activations:
$z(x)=z^{enc}(x)\cdot z^{*}(x)$ (2)
where $z^{*}(x)=f(A_{2}x+B_{2})$ has $A_{2}$ and $B_{2}$ that represent the weight matrix and bias. A visualization of the proposed architecture can be found in Fig. 3.
Vanishing Gradients
@@
Our key hypothesis is that the Latent assistance Network can prevent neurons to saturate easily, hence mitigating vanishing gradients. To explain this phenomenon, we investigate the derivative of a product of two functions. For the product of two arbitrary functions $g(x)\cdot h(x)$, the derivative is defined as $g^{\prime}(x)h(x)+g(x)h^{\prime}(x)$. In the context of using a sigmoid activation function for $f(x)$, the derivative of $z(x)$ becomes:
$z^{\prime}(x)=A_{1}\sigma(A_{1}x+B_{1})(1-\sigma(A_{1}x+B_{1}))\sigma(A_{2}x+B_{2})$
$+A_{2}\sigma(A_{1}x+B_{1})\sigma(A_{2}x+B_{2})(1-\sigma(A_{2}x+B_{2}))$
@@
$h^{\prime}(x)=A_{1}\text{sech}^{2}(A_{1}x+B_{1})\tanh(A_{2}x+B_{2})+\tanh(A_{1}x+B_{1})A_{2}\text{sech}^{2}(A_{2}x+B_{2})$
where $\operatorname{sech}^{2}$ is the derivate of a hyperbolic tangent. When using a hyperbolic tangent, the key difference with a sigmoid is that for any saturating value of $g(x)$, it will consistently keep the second part of the equation nonzero and dominated by $h(x)$, due to the hyperbolic tangents’ saturation into strictly nonzero values. This means that during saturation of a neuron in $g(x)$, the non-saturated function $h(x)$ can still maintain a non-trivial gradient in the product derivative $h^{\prime}(x)$, thus providing a mechanism to avoid vanishing gradients.
Taking a more formal approximation of neuron collapse through the lens of probability theory, we define the probability of a single neuron saturating as $\mathbf{p}$. Furthermore, in the case of a sigmoid or hyperbolic tangent, we assume symmetric saturation probabilities to both ends, defining the probability of a neuron saturating to one end of the spectrum as $0.5\mathbf{p}$. Under these assumptions, we show that interpreting a neuron as the product of two individual neurons will change saturation probabilities depending on the neuron activation function. For an overview, we refer the reader to both Table 1 and the corresponding empirical evidence in Appendix C.2.
Hyperbolic Tangent
@@
ReLU
In the case of a ReLU activation, we also assume the probability of a single neuron dying to be $\mathbf{p}$. As we look at the product of two neurons, the probability that one of the two neurons does not saturate is therefore $1-\mathbf{p}$, and the probability that both neurons do not saturate is $(1-\mathbf{p})^{2}$. The probability that at least one neuron saturates is thus equal to $1-(1-\mathbf{p})^{2}=2\mathbf{p}-\mathbf{p}^{2}$. As the ReLU saturation results in strict zeroes, this results in the product also being zero. Taking a product of ReLU activated neurons therefore increases the final neuron saturation probability from $\mathbf{p}$ to $2\mathbf{p}-\mathbf{p}^{2}$.
Table 1: Summary of dying neuron probabilities with and without a Latent assistance Network (LaN).
Activation Function Dying Neuron Probability Probability with LaN Difference
@@
It might seem logical to address neuron saturation by increasing the representation layer size, e.g. doubling it. However, in the experiments section we show that larger layers are not desirable and fail to achieve the unique benefits of the introduced architecture.
Furthermore, the product function $h(x)g(x)$ can exhibit more complex behavior due to the interactions between the input parameters influenced by both functions. These properties are illustrated with a basic regression task in Fig. 2.
(a) Tanh - $10^{6}$ iterations
(b) Tanh - $10^{7}$ iterations
(c) Tanh (LaN) - $10^{6}$ iterations
(d) Tanh (LaN) - $10^{7}$ iterations
Figure 4: Kernel Density Estimations (KDE) over a subset of 16 neurons in the compressed representation $z_{t}$ after training DQN in the ’Breakout’ environment. Each neuron represents one dimension of the representation $z_{t}\in\mathbb{R}^{512}$. The representation using a LaN tends to more quickly utilize the full range of the hyperbolic tangent and strongly mitigates dying neurons as compared to using a single hyperbolic tangent. Furthermore, using a LaN leads to more multimodal density distribution as compared to using a regular hyperbolic tangent. More information on the neuron KDE’s can be found in Appendix B.
Dying Hyperbolic Tangents
Although common literature has focused on the dying ReLU problem (He et al., 2015; Lu et al., 2019; Gulcehre et al., 2022; Sokar et al., 2023), we find that representations activated by hyperbolic tangents tend to also show strong dying neuron behavior. Given that the hyperbolic tangent is an asymptotic function near its saturation point, the classification of its saturation remains qualitative. In order to combine visualization and neuron classification, we approximate the amount of dying hyperbolic tangents by applying a Kernel Density Estimation (KDE) to the activations $\alpha_{i},i\in\mathbb{R}^{w}$ of each individual neuron in the compressed representation $z_{t}$. More details of the KDE calculation and neuron classification can be found in Appendix B. For the visualization of the representation layer, a fixed subset of the neurons $\alpha_{i}$ is taken. For both an encoder with and without using a LaN, we provide a visualization of sixteen individual neuron KDE’s in Fig. 4.

### v2

# Hadamard Representations: Augmenting Hyperbolic Tangents in RL

@@ Abstract
Activation functions are one of the key components of a deep neural network. The most commonly used activation functions can be classed into the category of continuously differentiable (e.g. tanh) and linear-unit functions (e.g. ReLU), both having their own strengths and drawbacks with respect to downstream performance and representation capacity through learning (e.g. measured by the number of dead neurons and the effective rank). In reinforcement learning, the performance of continuously differentiable activations often falls short as compared to linear-unit functions. We provide insights into the vanishing gradients associated with the former, and show that the dying neuron problem is not exclusive to ReLU’s. To alleviate vanishing gradients and the resulting dying neuron problem occurring with continuously differentiable activations, we propose a Hadamard representation. Using deep Q-networks and proximal policy optimization in the Atari domain, we show faster learning, a reduction in dead neurons and increased effective rank.
@@ 1 Introduction
Figure 1: Normalized performance training DQN in the Atari domain where the nonlinear activation function of the representation, defined as the final hidden layer, is either ReLU or tanh. A strong performance discrepancy due to a relatively small architectural change can be observed.
@@
Out of all activation functions, the Rectified Linear Unit (ReLU) (Fukushima, 1969; Nair & Hinton, 2010) and its variants (Xu et al., 2015; Klambauer et al., 2017) have emerged as the most widely used and generally best-performing activation functions up until this day (Jarrett et al., 2009; Goodfellow et al., 2016). The strength of the ReLU activation lies in its ability to naturally avoid vanishing gradients when used in deeper networks, in contrast to the continuously differentiable activation functions, such as the sigmoid and the hyperbolic tangent (Glorot & Bengio, 2010).
@@
A common drawback of using the ReLU activation is its limited expressivity in the context of shallow networks (see Fig. 2), as well as the phenomenon known as the dying ReLU problem (He et al., 2015; Lu et al., 2019). As training progresses, the number of dying ReLU’s tend to increase, resulting in a dying network and loss of network capacity (Dubey et al., 2022).
@@
In reinforcement learning (RL) (Sutton & Barto, 2018), even though training results in a large number of dying ReLU’s (Gulcehre et al., 2022; Sokar et al., 2023), the ReLU function still remains the most popular activation for performance reasons (Henderson et al., 2018). Similar to supervised learning (Teney et al., 2024), continuously differentiable activation functions such as the hyperbolic tangent are therefore not commonly used in RL (see Fig. 1). However, one might argue that their symmetrical, bounded shape and smooth gradient landscape offer optimization advantages that the ReLU lacks. Recent findings also indicate that a hidden layer activated by a hyperbolic tangent displays a high effective rank and thus a high layer expressivity (Kumar et al., 2021; Gulcehre et al., 2022). Despite being a theoretically sound candidate, its lack of success in RL has not been thoroughly investigated.
We therefore aim to provide insights into the hyperbolic tangent’s suboptimality, revealing that the vanishing gradient problem leads to dying neurons and under-utilization of the full network capacity. Based on these insights, we mitigate said effects by augmenting the original architecture. Specifically, we focus our research on the activations of the compressed representation of an encoder, defined as the final hidden layer, and provide an alternative to the conventional parameterization of this representation. Our contributions can be summarized as follows:
@@
  • We show that, in reinforcement learning, dying hyperbolic tangents are a phenomenon of a similar scale as the dying ReLU problem.
  • A Hadamard representation (HR) is proposed, defining a latent representation as the Hadamard product of two separate, individually parameterized activation vectors.
@@
  • We empirically show that, without hyperparameter tuning or the use of auxiliary losses, a Hadamard representation yields significant performance gains over a standard hyperbolic tangent in the Atari domain, and reveal how it decreases dying neurons and increases the internal representations’ effective rank.
@@ 2 Preliminaries
Figure 2: A regression of three shallow neural network architectures on a random complex sinusoidal function ($y=10*torch.sin(7*x)+15*torch.sin(10*x)+5*torch.cos(5*x)$). The Tanh (HR) network emerges as the strongest function approximator, even while having less trainable parameters (501 vs 601 for Tanh & ReLU). To make a fair comparison, the Tanh and ReLU networks have one single hidden layer of size 200, while the Tanh (HR) network has a hidden layer of size 100. For the Tanh (HR) network however, we use two parallel linear layers preceding the hidden layer in order to be able to use the single hidden layer as the Hadamard product of two activations (see Section 3). For experiments comparing deeper networks, we refer the reader to Appendix C.1.
We consider an agent acting within its environment, where the environment is modeled as a discrete Markov Decision Process (MDP) defined as a tuple $(\mathcal{S},\mathcal{A},T,R,\gamma)$. $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the environment’s transition function, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{R}$ is the environment’s reward function and $\gamma\in[0,1)$ is the discount factor. A replay buffer $B$ is used to store visited states $s_{t}\in\mathcal{S}$ that were followed by actions $a_{t}\in\mathcal{A}$ and resulted in the rewards $r_{t}\in\mathcal{R}$ and the next states $s_{t+1}$. One entry in $B$ contains a tuple of past experience $(s_{t},a_{t},r_{t},s_{t+1})$. The agent’s goal is to learn a policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ that maximizes the expectation of the discounted return $V^{\pi}(s)=\operatorname{\mathbb{E}}_{\tau}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\mid s_{t}=s]$, where $\tau$ is a trajectory following the policy $\pi$.
@@
3 Augmenting Hyperbolic Tangents
@@
Dying Hyperbolic Tangents
Although common literature has focused on the dying ReLU problem (He et al., 2015; Lu et al., 2019; Gulcehre et al., 2022; Sokar et al., 2023), we find that hidden layers activated by hyperbolic tangents similarly show strong dying neuron behavior. When using a hyperbolic tangent in a deep neural network, a single neuron $\alpha_{i},i\in\mathbb{R}^{w}$, with $w$ the layer dimension, can be classified as dying if:
$|\alpha_{i}|\approx 1\;\forall\;s_{t}\in B$ (2)
Where $s_{t}$ is an observation in buffer $B$. Note that an absolute value is used, as a hyperbolic tangent can either saturate towards -1 or 1. Given that the hyperbolic tangent is an asymptotic function near its saturation point, an approximate equality must be considered, as the classification of its saturation will always remain qualitative ($|\alpha_{i}|\neq 1\;\forall\;s_{t}\in B$) . To approximate the condition given in Eq. 2, the amount of dying hyperbolic tangents is calculated by using a kernel density estimation (KDE) (Silverman, 1986) on the activations $\alpha_{i},i\in\mathbb{R}^{w}$ of each individual neuron in the activation layer. In order to visualize activations in a hidden layer, a fixed subset of the KDE’s of the neurons $\alpha_{i}$ is taken. A clear visualization of dying hyperbolic tangents during training in the Atari Breakout environment can be seen by analyzing sixteen individual neuron KDE’s in Fig. 3.
(a) Tanh - $10^{6}$ iterations
(b) Tanh - $10^{7}$ iterations
Figure 3: Kernel Density Estimations (KDE) over a subset of 16 neurons in the compressed representation $z_{t}$ after training DQN Mnih et al. (2015) in the Breakout environment using a hyperbolic tangent activation for $z_{t}$. Each neuron represents one dimension of the representation $z_{t}\in\mathbb{R}^{512}$. Red outlines represent dying neurons, where a near infinite sized density spike occurs at either 1 or -1.
More details on the KDE calculation and dying neuron classification can be found in Appendix B.
Hadamard Representations (HR)
@@
As Fig. 3 indicates that the activation of $z_{t}$ with a hyperbolic tangent leads to saturation and dying neurons, an augmentation of the representation architecture is proposed. In the conventional encoder setting, the networks’ final hidden layer is defined as $z^{enc}(x)=f(A_{1}x+B_{1})$, with $A_{i}$ and $B_{i}$ representing weight and bias parameters, the function $f()$ representing a nonlinear activation function while $x$ is the set of activations from the previous layer. In order to reduce the information dependence on a single set of neurons, we propose using a Hadamard representation that augments the original representation with a parallel representation layer $z^{*}$. This can be interpreted as using a single highway layer with a closed carry gate (Srivastava et al., 2015). The final representation is defined as the Hadamard product between the aforementioned activations $z(x)=z^{enc}(x)\cdot z^{*}(x)$, where $z^{*}(x)=f(A_{2}x+B_{2})$. A visualization of the proposed architecture can be found in Fig. 4.
Figure 4: A visualisation of the Hadamard Representation architecture combined with the nature DQN architecture applied on a snapshot of the ‘SpaceInvaders’ Atari environment (Mnih et al., 2015). Vertical and horizontal bars represent convolutional and linear layers, respectively. A parallel linear layer providing independently parameterized hidden layer activations is integrated, where the Hadamard product of the two activations represents the final latent representation $z_{t}$ used for downstream learning.
Preventing Dying Neurons
@@
Our key hypothesis is that the Hadamard representation can prevent saturation, hence alleviating vanishing gradients and dying neurons (See Eq. 2). To support our hypothesis, we investigate the derivative of a product of two functions. For the product of two arbitrary functions $g(x)\cdot h(x)$, the derivative is defined as $\uline{\color[rgb]{0.875,0.26171875,0.8984375}{\color[rgb]{0,0,0}g^{\prime}(x)h(x)}}+\uline{\color[rgb]{0.296875,0.76171875,0.83203125}{\color[rgb]{0,0,0}g(x)h^{\prime}(x)}}$. In the context of using a sigmoid activation function for $f(x)$, the derivative of $z(x)$ becomes:
$z^{\prime}(x)=\uline{\color[rgb]{0.875,0.26171875,0.8984375}{\color[rgb]{0,0,0}A_{1}\sigma(A_{1}x+B_{1})(1-\sigma(A_{1}x+B_{1}))\sigma(A_{2}x+B_{2})}}\color[rgb]{0,0,0}$
$+\uline{\color[rgb]{0.296875,0.76171875,0.83203125}{\color[rgb]{0,0,0}A_{2}\sigma(A_{1}x+B_{1})\sigma(A_{2}x+B_{2})(1-\sigma(A_{2}x+B_{2}))}}$
@@
$h^{\prime}(x)=\uline{\color[rgb]{0.875,0.26171875,0.8984375}{\color[rgb]{0,0,0}A_{1}\text{sech}^{2}(A_{1}x+B_{1})\tanh(A_{2}x+B_{2})}}\color[rgb]{0,0,0}+\uline{\color[rgb]{0.296875,0.76171875,0.83203125}{\color[rgb]{0,0,0}A_{2}\text{sech}^{2}(A_{2}x+B_{2})\tanh(A_{1}x+B_{1})}}$
where $\operatorname{sech}^{2}$ is the derivative of a hyperbolic tangent. When using a hyperbolic tangent, the key difference with a sigmoid is that for any saturating value of $g(x)$, it will consistently keep the second part of the equation nonzero and dominated by $h(x)$, due to the hyperbolic tangents’ saturation into strictly nonzero values. This means that during saturation of a neuron in $g(x)$, the non-saturated function $h(x)$ is still able to maintain a non-trivial gradient in the product derivative $h^{\prime}(x)$, thus providing a mechanism to avoid vanishing gradients. We visualize the kernel densities during training with a Hadamard representation in Fig. 5. The individual representations before taking the Hadamard product can be found in Appendix B.
Taking a more formal approximation of neuron collapse, we start by defining the probability of a single neuron saturating as $\mathbf{p}$. Furthermore, in the case of a sigmoid or hyperbolic tangent, we assume symmetric saturation probabilities to both ends, defining the probability of a neuron saturating to one end of the spectrum as $0.5\mathbf{p}$. Under these assumptions, we show that interpreting a neuron as the product of two individual neurons will change saturation probabilities depending on the neuron’s activation function.
(a) Tanh (HR) - $10^{6}$ iterations
(b) Tanh (HR) - $10^{7}$ iterations
Figure 5: Kernel Density Estimations (KDE) over a subset of 16 neurons in the compressed representation $z_{t}$ after training DQN in the Breakout environment using a Hadamard representation (HR) with hyperbolic tangents. The Hadamard representation tends to quickly utilize the full range of the hyperbolic tangent while also mitigating dying neurons.
Hyperbolic Tangent:
@@
Rectified Linear-Unit:
In the case of a ReLU activation, we also assume the probability of a single neuron dying to be $\mathbf{p}$. As we look at the product of two neurons, the probability that one of the two neurons does not saturate is therefore $1-\mathbf{p}$, and the probability that both neurons do not saturate is $(1-\mathbf{p})^{2}$. The probability that at least one neuron saturates is thus equal to $1-(1-\mathbf{p})^{2}=2\mathbf{p}-\mathbf{p}^{2}$. As the ReLU saturation results in strict zeroes, this results in the product also being zero. Taking a product of ReLU activated neurons therefore increases the final neuron saturation probability from $\mathbf{p}$ to $2\mathbf{p}-\mathbf{p}^{2}$. For an overview, we refer the reader to both Table 1 and the corresponding empirical evidence in Appendix C.3.
Table 1: Predicted dying neuron probabilities with and without a Hadamard representation.
Activation Function Dying Neuron Probability Probability with HR Difference